# UINav: A maker of UI automation agents

### v1

@@ Abstract
An automation system that can execute natural language instructions by driving the user interface (UI) of an application can benefit users, especially when situationally or permanently impaired. Traditional automation systems (manual scripting, programming by demonstration tools, etc.) do not produce generalizable models that can tolerate changes in the UI or task workflow. Machine-learned automation agents generalize better, but either work only in simple, hand-crafted applications or rely on large pre-trained models, which may be too computationally expensive to run on mobile devices. In this paper, we propose UINav, a demonstration-based agent maker system. UINav agents are lightweight enough to run on mobile devices, yet they achieve high success rates with a modest number of task demonstrations. To minimize the number of task demonstrations, UINav includes a referee model that allows users to receive immediate feedback on tasks where the agent is failing to best guide efforts to collect additional demonstrations. Further, UINav adopts macro actions to reduce an agent’s state space, and augments human demonstrations to increase the diversity of training data. Our evaluation demonstrates that with an average of 10 demonstrations per task UINav can achieve an accuracy of 70% or higher, and that with enough demonstrations it can achieve near-perfect success rates on 40+ different tasks.
@@ 1 Introduction
The next frontier in artificial intelligence are agents that can autonomously operate computers as humans do. Instructed by users in natural language, these agents are especially valuable when their users have visual or motor disabilities, when they are situationally impaired (e.g., driving, cooking), or when they are engaged in learning new software features and can benefit from in-context guidance. Towards this goal, recently, we have seen a growing interest in agents that fulfill human tasks by manipulating UI elements on a screen Adept (2023), by invoking service APIs Nakano et al. (2021); Komeili et al. (2022), or by interacting with software tools Parisi et al. (2022); Schick et al. (2023). Our work falls in the first category, UI automation, where we propose a demonstration-based agent building system designed to produce lightweight agents that can execute tasks specified in natural language with near perfect success rates.
@@
Existing approaches to UI automation range from manual scripting to large Transformer-based models. Manually-coded UI scripts can work reliably, but they are expensive to write and maintain. This is what enterprises currently use to automate their workflows UIPath (2023); Microsoft Power Automate (2023); Automation Anywhere (2023). Programming by demonstration systems provide a more natural interface, where a user can record a sequence of UI interactions and replay them when needed Kundra (2020); Barman et al. (2016); Riva and Kace (2021); Li et al. (2017). However, these systems do not produce generalizable models that can tolerate changes in user interfaces and workflows, thus leading to high maintenance costs. On the other hand, machine learning approaches to UI automation aim to scale to many tasks and applications. Some previous approaches use imitation learning and reinforcement learning Liu et al. (2018); Gur et al. (2018). However, their testing environments Shi et al. (2017) are much simpler than real-world websites. Moreover, their reward model either needs to be manually programmed or requires large numbers of demonstrations Li and Riva (2021); Humphreys et al. (2022). More recent work Li et al. (2020a); Li and Li (2023); Venkatesh et al. (2022); Wang et al. (2023) leverages the power of Transformers and pre-trained large language models (LLMs) such as PaLM Chowdhery et al. (2022) and GPT-3 Brown et al. (2020). Despite the improvement in performance, these models incur high resource costs (multiple days of training on hundreds of GPUs/TPUs and inference costs beyond the capability of modern devices) which may not always be affordable.
@@
Trade-offs are required between accuracy, generalizability and computational costs when selecting an approach to UI automation. We aim to find a sweet spot between these three properties. We propose UINav, an agent maker system designed to produce lightweight neural agents that can run on mobile devices while yielding high success rates. With UINav, users record task demonstrations and quickly train neural agents to test in the wild.
@@
As in prior work, UINav needs to address the fundamental challenge of how to achieve acceptable success rates with a small number of demonstrations. We observe that the number of demonstrations required to achieve good performance differs widely across tasks and execution environments. If the environment is relatively static even a handful of demonstrations is sufficient to learn. On the other hand, for tasks that have to work across many different UIs more demonstrations may be needed. UINav allows users to quickly and continuously evaluate the performance of intermediate versions of their agents with new tasks and in new environments until a desired quality is achieved. They can do so by using the UINav referee model. This model is trained with the same set of demonstrations used to train the automation agent, but with a different goal, i.e., predicting whether a task is successfully completed (rather than predicting which UI elements to interact with). When testing their agents, users receive immediate feedback on which tasks are failing and may benefit from additional demonstrations.
@@
Another challenge UINav addresses is how to increase the robustness of automation agents against system delays and changes in the UI. It does so through two key techniques. First, every UI action is executed as a small program composed of lower level operations with status checks. These programs, referred to as macro actions, abstract the system-specific details thus greatly reducing the agent’s state space and therefore the number of required demonstrations. Second, UINav adopts demonstration augmentation where human demonstrations are augmented by randomizing non-critical UI elements to increase their diversity.
We develop UINav using an internal dataset of 40+ tasks. We also evaluate it on a public dataset, where UINav outperforms state-of-the-art systems by large margins. Overall, this work makes the following contributions:
  • A practical system designed to build UI automation agents from scratch that achieve near perfect success rates on previously seen tasks and that can be deployed to mobile devices.
  • An error-driven process to collect demonstrations paired with augmentation techniques and macro actions which overall minimize the number of required human demonstrations and increase robustness.
  • A comprehensive evaluation demonstrating UINav’s advantages over state-of-the-art systems.
@@ 2 Related work
2.1 UI automation scripts
UI automation scripts consist of low-level action statements (click, type, enter, etc.) each referring to a specific UI element (button, input field, image, etc.). Record-and-replay tools like Selenium Kundra (2020) can be used to facilitate the generation of automation scripts. These scripts can also be wired together and interleaved with API calls to obtain full automation workflows, using robotic process automation tools UIPath (2023); Automation Anywhere (2023); Blue Prism (2023).
@@ 2.1 UI automation scripts
To make automation accessible also to non-technical users, programming by demonstration tools have been proposed Sugiura and Koseki (1998); Leshed et al. (2008); Lin et al. (2009); Li et al. (2010); Barman et al. (2016); Li et al. (2017); Chasins et al. (2018). These are advanced record-and-replay tools that can generate fully functional automation scripts and even action graphs Riva and Kace (2021) from recordings of task interactions (demonstrations). Some work also explores whether these demonstrations can be provided in the format of video recordings Bernal-Cárdenas et al. (2020); Chen et al. (2022). Overall, a major downside of this line of work is that these systems do not produce models that generalize to new task workflows and UIs. Every new task and every variation to an existing task requires new demonstrations or manual updates.
To limit the collection of new demonstrations but still provide useful automation, other work examines this problem from an interaction standpoint. PUMICE Li et al. (2019) enables users to resolve ambiguities in their natural language instructions by collaborating with the agent by introducing new concepts and learning which actions the agent supports. SOVITE Li et al. (2020b) displays the system’s understanding of user intents using UI screenshots, thus allowing users to identify failures of the system in understanding their instructions and to correct them using direct manipulation on the screenshots. Unfortunately, these approaches incur an overhead on the user side and make the interaction with the agent less fluid. Moreover, their ability of learning from prior user feedback is limited or completely absent.
2.2 ML-based automation
@@
ML-based automation systems promise to reduce the overhead of UI automation scripts by abstracting single traces into general models. Transformer-based architectures Li et al. (2020a); Bai et al. (2021); He et al. (2021); Banerjee et al. (2022); Li and Li (2023) and reinforcement learning approaches Liu et al. (2018); Gur et al. (2018); Li and Riva (2021) have been proposed to train agents capable of navigating apps and websites when provided with natural language instructions. Yet, it is unclear how well these systems perform in a variety of real-world environments and scale across task categories. Reinforcement learning agents have been tested in synthetic webpages containing only 10–50 UI elements Shi et al. (2017). Existing datasets are mostly for Android: PixelHelp Li et al. (2020a) covers 4 apps for a total of 187 instructions; MoTIF Burns et al. (2022), which we use in this paper, consists of 6,100 episodes across 125 Android apps. As these datasets are collected using app crawling infrastructures and costly human labeling, they make large-scale training difficult.
Recent work has sought to leverage large language models such as PaLM and GPT3 to map natural language instructions into step-by-step instructions (parsing phase) which can be executed in the UI by a simpler grounding model (grounding phase). An example of such a system is UGIF Venkatesh et al. (2022). Wang et al. Wang et al. (2023) also prompt LLMs to perform conversations with a mobile UIs. A main downside of these approaches is the large training overhead (e.g., multiple days of training on hundreds of GPUs/TPUs) as well as the cost of running inference, which is in general too expensive to run on mobile devices.
@@
In this paper, we extend our previous work Li (2021) where macro actions were introduced but was limited to work with OCR and icon recognition, into a full system, that bridges the gap between programming by demonstrations and ML systems by providing an easy-to-learn system to train robust, multi-task AI agents for UI navigation in real-world applications. While the system requires manual demonstrations for training, it provides an error-driven collection of demonstrations where testing scenarios are automatically generated and evaluated by the system, thus reducing the number of redundant demonstrations. The error driven demo collection of UINav is inspired by the DAGGER Ross et al. (2011) algorithm and we show that it is effective in reducing the number of demonstrations for both sequential (UINav referee) and non-sequential (UINav agent) models.
In the next sections we introduce the UINav system. We start by providing some concrete examples of the challenges the system must face with motivates its focus on (i) reducing the number of user demonstrations but also (ii) making demonstrations as diversified as possible.
3 Challenges in UI automation
@@
In this paper we study how UI automation systems can generalize to new execution environments, including different apps and different states in the app, without requiring an excessive number of demonstrations. To illustrate the challenges involved we use an apparently simple task, search, i.e., operating the search bar of an application. The task typically involves typing some search terms into a text box and clicking a search button or pressing Enter.
@@
Two aspects make this task challenging. First, search is a universal task that must work across a large variety of apps where the search bar can take different formats. Some search bars require the user to click on a search icon (typically on the right hand-side) after typing a search term, while others display search results as the user types in; typing in a search bar can automatically open a drop-down menu with autocomplete suggestions or matching products with pictures and names; other search bars have an additional field (e.g., category) typically displayed on the left side of the search bar which must be set beforehand; there are also search bar that are hidden and reveal only after the user clicks on the search icon; etc.
@@
The second axis of complexity regards the start state. When the user initiates a search, the environment may be displaying a page with no search functionality or it may be in an app that is not the target of the search command. In these cases, the first problem the agent has to solve is how to navigate to the state offering the search function, which may involve navigating back in the current app, launching a different app, dismissing welcome screens or ads, etc. Even when the environment already displays the desired search widget, its state may need to be reset, by deleting search terms entered in a previous search or changing the search category. Figure 1 shows the UINav agent searching in YouTube. The agent dismisses popups twice (a) and (b) to reveal the search bar. It then clicks the X button to erase the previous search phrase “something” (c). The system does not reach the desired start state for a search until the screen shown in (d), where the agent sets the focus on the search bar to then enter the search term.
Figure 1: UINav agent searches in YouTube. The pink arrows highlight agent’s actions that are annotated by red boxes and texts. (a) Clicks the back button to dismiss a popup ads; (b) Clicks "X" to dismiss the install page of Audible; (c) Clicks "X" to erase the previously entered search phrase “something”; (d) Focus on the search bar to enter a new search term.
In general, in a real execution environment, an agent is exposed to many different screen conditions caused by a combination of factors, such as different apps, dynamic contents of the apps, previous interactions, layout variance on different devices, UI changes across different app versions and operating systems, ads and notifications, etc. An agent needs to have the capability of ignoring irrelevant UI elements and navigating to relevant states.
One way to model this variability is through more demonstrations, but with a clear overhead. UINav’s first contribution is to adopt an error-driven approach to reduce the number of required demonstrations (§4.2). Its second contribution is to amplify the learning brought by each demonstration by automatically augmenting them (§5.2). Finally, to addresses variability issues due to system delays rather than relying on demonstrations, UINav take a programmatic approach by introducing macro actions (§5.1).
@@ 4 System design
Figure 2: Agent and referee development loop and UINav high-level architecture.
Figure 2 illustrates the development workflow and high level architecture of UINav. Given a collection of tasks, each represented by a natural language utterance (e.g., "Add Max Smith to my contacts") and an observation of the current state (e.g., a representation of what is currently displayed on the device’s screen such as the home page or an application page), a neural network-backed agent responds with its choice of action to complete the task. The predicted action is executed by the environment by directly interacting with a device’s system (an emulator or a real phone). Then, the agent is provided with a new observation describing the new state and a new action is predicted. This setup is similar to that of standard reinforcement learning agents, but in UINav the environment also includes a second agent called referee, which is also backed by a neural network. The referee is responsible for judging the completion status of a task (also called episode) at each time step.
In developing UINav agents (left of Figure 2), first, some human demonstrations completing the target tasks are collected. Then, the neural networks of the agent (§4.1.2) and referee (§4.2) are trained (Both networks are initialized with random weights). Next, the trained agent and referee are evaluated on the device system. Failures of either the agent or the referee are recorded and used to guide the collection of new demonstrations which can be used in the next round of training. The development loops over these steps until no more errors of either the agent or the referee are found.
@@ 4.1 Agent’s neural network architecture
We now describe the architecture of the agent model (Figure 3). We focus the discussion particularly on the techniques that enable this architecture to reduce the number of required human demonstrations.
4.1.1 Screen representation
@@
An agent receives observations describing the current state of the device, which is represented by the screen currently displayed. A screen consists of a collection of UI elements, with each element described by a set of attributes: type (button, icon, etc.), text description (visible text, content description, resource identifier, etc.), position, utterance matching (whether on-screen text matches words in the utterance, see §4.3), and state (e.g., whether a checkbox is selected). The screen representation can be generated from raw pixels processed by screen understanding techniques Chen et al. (2020); Wu et al. (2021); Zhang et al. (2021), which also include icon detection and text recognition, or from a tree-structured representation of the UI, such as the Android Accessibility View Hierarchy. Our implementation dynamically switches between the two sources of screen representation based on simple heuristics, such as whether the target app is known to provide poor accessibility support or whether the number of accessibility nodes is extremely small (see §7 for more details).
Figure 3: The neural network architecture of the agent model.
4.1.2 Agent model
@@
Figure 3 shows the architecture of the agent model. The input to the neural network is a set of UI elements and an utterance. Each UI element is represented by a vector concatenated from the feature vectors of its attributes. Text labels of UI elements are encoded by a language model such as BERT Devlin et al. (2019). The feature vectors of the UI elements are fed into a Transformer encoder. The output of the Transformer encoder is a function of the encoding of each UI element plus its attention over all other UI elements on the screen, including itself. The utterance is encoded by the language model after its variable sub-strings are masked out (see §4.3 for details).
@@
Given these inputs the agent predicts which action to perform next. This involves predicting: the UI element on which to perform the action, the type of action, and any argument for the action. Our decoder uses a single cross-attention module, with the embedding of the masked utterance serving as the query vector and element encodings from the encoder serving as keys and values. The largest attention weight is used to select the element to act upon, while the vector output of the cross attention module is passed through two independent multi-layer perceptrons (MLP) to predict action type and argument. Together, the encoder and decoder represent a lightweight Transformer model.
@@
In its essence, the agent’s neural network is a scoring system. For any given screen, all the UI elements are scored, and the one with the highest score is selected. Due to the attention in the encoder, for any UI element, its relationship with all the other UI elements on screen can be encoded. Intuitively, the transformer model learns how different combinations of UI elements, and commands map to actions, and uses this knowledge to rank elements to act on. For UI screens it is particularly important that the model learns to evaluate single UI elements in the context of others because the meaning of UI elements is often context sensitive Banerjee et al. (2022) – elements with similar appearance (color, size and shape) can have very different functions but neighboring elements such as text labels or other UI controls can help resolve the ambiguity. For more evidence on how UINav effectively learns to contextually evaluate UI elements, please find some examples in §A.1.2 and §A.1.3.
4.1.3 Action space
Table 1: UINav action space.
Element click <elem>
Clicks the center of the element.
actions focus_and_type <elem,text>
Sets focus on the element, types the text, and optionally presses Enter.
dismiss <elem>
Clicks outside of the element.
Global wait
Waits until the next observation.
actions back
Goes back to the previous screen.
scroll <left|right|up|down>
Scrolls in the specified direction.
open_app <app_name>
Launches the specified application.

### v2

# UINav: A Practical Approach to Train On-Device Automation Agents

@@ Abstract
Automation systems that can autonomously drive application user interfaces to complete user tasks are of great benefit, especially when users are situationally or permanently impaired. Prior automation systems do not produce generalizable models while AI-based automation agents work reliably only in simple, hand-crafted applications or incur high computation costs. We propose UINav, a demonstration-based approach to train automation agents that fit mobile devices, yet achieving high success rates with modest numbers of demonstrations. To reduce the demonstration overhead, UINav uses a referee model that provides users with immediate feedback on tasks where the agent fails, and automatically augments human demonstrations to increase diversity in training data. Our evaluation shows that with only 10 demonstrations UINav can achieve 70% accuracy, and that with enough demonstrations it can surpass 90% accuracy.
UINav: A Practical Approach to Train On-Device Automation Agents
Wei Li, Fu-Lin Hsu, Will Bishop, Folawiyo Campbell-Ajala, Max Lin, Oriana Riva Google Research
@@ 1 Introduction
The next frontier in artificial intelligence is agents that autonomously operate computers as humans do. Instructed by users in natural language, these agents are especially valuable when their users have visual or motor disabilities or when they are situationally impaired (e.g., driving, cooking). We are particularly interested in agents that can execute human tasks by interacting directly with the user interface (UI) of a running application. These so-called UI automation agents Liu et al. (2018); Li et al. (2020); Humphreys et al. (2022) can scale well to support a myriad of tasks because they do not depend on third-party APIs.
@@
Existing approaches to UI automation range from UI scripting to AI-based agents. UI scripts can work reliably, but they involve coding or manual demonstrations Kundra (2020); Barman et al. (2016); Riva and Kace (2021); Li et al. (2017) and they cannot tolerate well changes in the UI and workflows, thus leading to high maintenance costs – this is, however, what enterprises use to automate business workflows UIPath (2023). AI-based approaches can scale better. Using imitation learning and reinforcement learning Liu et al. (2018); Gur et al. (2018), agents are trained to navigate the web autonomously. However, their synthetic and simplified test environments Shi et al. (2017) and their dependency on large amounts of demonstrations Humphreys et al. (2022) make them hard to deploy. Recent work leverages Transformers Li et al. (2020); Li and Li (2023); Venkatesh et al. (2022); Wang et al. (2023) and pre-trained large language models (LLMs) Yan et al. (2023); Venkatesh et al. (2022); Zheng et al. (2024). Despite the performance improvement, these solutions come with large resource costs (multiple days of training on hundreds of GPUs/TPUs and high inference costs).
@@
A practical approach to UI automation requires trading between accuracy, generalizability and computational costs. We find a sweet spot between these three properties, and propose UINav, a demonstration-based system designed to produce lightweight neural agents that can run on mobile devices while yielding good success rates.
@@
As in prior work, UINav needs to address the challenge of how to achieve good success rates with fewer demonstrations. We observe that the demonstrations required to achieve good performance differs widely across tasks and environments. If the environment is relatively static even a handful of demonstrations is sufficient; for tasks that must work across many different UIs more demonstrations are needed. When collecting demonstrations, UINav provide users with immediate feedback on which tasks are failing and may benefit from additional demonstrations, and which are satisfactory. It does so through a referee model which is trained with the same set of demonstrations used to train the automation agent, but with a different goal: predicting whether a task is successfully completed (rather than predicting which UI action to perform).
@@
Another challenge UINav addresses is how to increase the robustness of automation agents against system delays and changes in the UI. It does so through three key techniques. First, every UI action is executed as a small program composed of lower-level operations with status checks. These programs, referred to as macro actions, abstract the system-specific details thus greatly reducing the agent’s state space and therefore the number of required demonstrations. Second, UINav adopts demonstration augmentation where human demonstrations are augmented by randomizing non-critical UI elements to increase their diversity. Finally, through utterance masking variable sub-strings in utterances are abstracted out.
We develop UINav using an internal dataset of 40+ tasks and test is on actual Android phones. We also evaluate it on a public dataset, where UINav outperforms various baselines and demonstrates generalizability. Overall, we make the following contributions: (i) a practical system to build UI automation agents that achieve near perfect success rates on previously seen tasks and that can be deployed to mobile devices; (ii) an error-driven process to collect demonstrations paired with augmentation techniques and macro actions; and (iii) a comprehensive evaluation demonstrating UINav’s advantages over state-of-the-art systems.
@@ 2 Related work
UI automation scripts.
@@
Record-and-replay tools like Selenium Kundra (2020) can be used to facilitate the generation of UI automation scripts. These scripts can also be integrated with robotic process automation tools UIPath (2023); Automation Anywhere (2023); Blue Prism (2023). Programming by demonstration tools Sugiura and Koseki (1998); Leshed et al. (2008); Lin et al. (2009); Li et al. (2010); Barman et al. (2016); Li et al. (2017); Chasins et al. (2018) are advanced record-and-replay tools that can generate fully functional UI scripts and even action graphs Riva and Kace (2021) from recordings of task interactions (demonstrations), which could also be provided in the format of video recordings Bernal-Cárdenas et al. (2020); Chen et al. (2022). Overall, a major downside of this line of work is that these systems do not produce models that generalize to new task workflows and UIs.
AI-based automation.
@@
Transformer-based architectures Li et al. (2020); Bai et al. (2021); He et al. (2021); Banerjee et al. (2022); Li and Li (2023) and reinforcement learning approaches Liu et al. (2018); Gur et al. (2018); Li and Riva (2021) have been proposed to train agents capable of navigating apps and websites when provided with natural language instructions. Yet, it is unclear how well these systems perform in a variety of real-world environments and scale across task categories because either they have been tested in synthetic webpages of 10–50 UI elements Shi et al. (2017) or on limited datasets Li et al. (2020); Burns et al. (2022). Recent work leverages LLMs to ground natural language instructions in UIs Venkatesh et al. (2022); Wang et al. (2023); Yan et al. (2023); Zheng et al. (2024); Rawles et al. (2023). These approaches come with a large training overhead (e.g., multiple days of training on hundreds of GPUs/TPUs) and a high inference cost which prevents them from running on mobile devices.
@@
In this paper, we extend our previous work Li (2021) where macro actions were introduced but was limited to work with OCR and icon recognition, into a full system, that bridges the gap between programming by demonstrations and AI-based systems by providing an easy-to-learn system to train robust, multi-task agents for UI navigation in real-world applications. While the system requires manual demonstrations for training, it provides an error-driven collection of demonstrations where testing scenarios are automatically generated and evaluated by the system, thus reducing the number of redundant demonstrations. The error driven demo collection of UINav is inspired by the DAGGER Ross et al. (2011) algorithm and we show that it is effective in reducing the number of demonstrations for both sequential (referee) and non-sequential (agent) models.
3 Why is UI automation hard?
@@
We study the problem of how a UI automation system can generalize to new execution environments, including different apps and different tasks, without requiring an excessive number of demonstrations. To illustrate the challenges we use an apparently simple task, search, i.e., operating the search bar of an app. Two aspects make this task challenging.
@@
Search is a universal task that must work across a myriad of apps where search bars can take many different formats. Some search bars require the user to type some keywords and then click an icon (typically on the right hand-side); others, as the user types, automatically display search results which can be directly opened; some others have an additional field (e.g., a category) which must be set beforehand; there are also search bars that are hidden and reveal only upon clicking on an icon; etc.
@@
The second axis of complexity regards the agent’s start state. When an agent is requested to search in a specific app, the user’s device screen may not display the target app or may display it in a page (state) without any search functionality. The agent must first understand how to navigate to the state offering the search function, which may involve navigating back, launching a different app, or dismissing welcome screens and ads. Even when the environment already shows the desired search widget, its state may need to be reset, e.g., by deleting search terms previously entered (see the YouTube example in Fig. 6 in the Appendix).
In general, in a real environment, an agent is exposed to many different screen conditions caused by a combination of factors: different apps, dynamic app content, previous interactions, layout variance across devices, UI changes across app/OS versions, ads and notifications, etc. An agent needs to ignore irrelevant UI elements and navigate to relevant states. One way to tackle this variability is through more demonstrations, but with obvious overheads. UINav’s first contribution is to adopt an error-driven process to collect demonstrations (§4.2). Its second contribution is to amplify the learning brought by each demonstration by automated augmentation (§5). Finally, to address variability issues due to system delays, rather than relying on demonstrations UINav takes a programmatic approach by introducing macro actions (§5).
@@ 4 System design
Figure 1: High-level architecture of UINav.
Fig. 1 shows the high-level architecture of UINav. Given a task represented by a natural language utterance and an observation of the device state (i.e., a representation of what is currently displayed on the screen), a neural network-backed agent responds with its choice of action to complete the task. The predicted action is executed by the environment by interacting with a device’s system (an emulator or a real phone). Then, the agent is provided with a new observation describing the new state and a new action is predicted. This setup is similar to that of a reinforcement learning agent, but UINav also includes a second agent called referee, which is responsible for judging the completion status of a task (episode) at each time step.
The development of UINav agents (left of Fig. 1) involves first collecting human demonstrations for some target tasks, then training the neural networks of the agent (§4.1) and referee (§4.2), and finally evaluating them on the device. Failures of either the agent or the referee are recorded and used to guide the collection of new demonstrations to be used in the next round of training. The development loops over these steps until no more errors of either the agent or the referee are found.
@@ 4.1 Agent’s neural network architecture
Figure 2: The neural network of the agent model.
@@
The UINav agent consists of an encoder-decoder architecture (Fig. 2). It perceives the state of the device through observations of what is currently displayed on the screen, represented by the set of UI elements composing it. Each UI element is described by a set of attributes: type (button, icon, etc.), text (visible text, content description, resource identifier, etc.), on-screen position, utterance matching (whether on-screen text matches the utterance [^1] Similarly to previous work Liu et al. (2018), we compute utterance matching as the average of the similarity scores of all words in the UI element’s text with the utterance.), and state (e.g., whether a checkbox is selected). The screen representation can be generated from raw pixels processed by screen understanding techniques Chen et al. (2020); Wu et al. (2021); Zhang et al. (2021), which also include icon detection and text recognition, or from a tree-structured representation of the UI, such as the Android Accessibility tree. Our implementation dynamically switches between the two sources of screen representation based on simple heuristics, such as whether the target app is known to provide poor accessibility support or whether the number of accessibility nodes is extremely small.
@@
Then, the input to the neural network of the agent is a set of UI elements and an utterance. Each UI element is represented by a vector concatenated from the feature vectors of its attributes. Text labels of UI elements are encoded by a language model Devlin et al. (2019). The feature vectors of the UI elements are fed into a Transformer encoder. The output of the encoder is a function of the encoding of each UI element plus its attention over all other UI elements on the screen, including itself.
The decoder predicts which action to perform. This involves predicting (i) the UI element on which to perform the action, (ii) the type of action (click, type, etc.), and (iii) any argument for the action. Actions (summarized in Table 3, §A.2) can be of two types. Element actions (click, focus_and_type, dismiss) manipulate a specific element, while global actions (wait, back, scroll, open_app) are general operations or platform-specific functions.
@@
The decoder uses a single cross-attention module, with the utterance embedding serving as the query vector and element encodings serving as keys and values. The largest attention weight is used to select the element to act upon, while the vector output of the cross-attention module is passed through two independent multi-layer perceptrons (MLP) to predict action type and argument.
@@
In its essence, the agent’s neural network implements a scoring system. For any given screen, all its elements are scored, and the highest-scored one is selected. Due to the attention in the encoder, for any UI element, its relationship with all the other elements can be encoded. The Transformer model learns how different combinations of UI elements and utterances map to actions, and uses this knowledge to rank elements to act on. It is essential that the model learns to evaluate single UI elements in the context of others because the meaning of UI elements is often context sensitive Banerjee et al. (2022) – elements of similar appearance (color, size and shape) can have different functions but neighboring elements like text labels can help resolve the ambiguity. For specific examples on how UINav contextually evaluates UI elements see §A.8.